OpenStudy (anonymous):

Twice the difference between a number and 4 is twenty. Whats the number?

OpenStudy (ankit042):

2x-4 = 20

OpenStudy (anonymous):

so 8 ... thanks (:

OpenStudy (johnweldon1993):

Not quite... 2(x - 4) = 20 Twice the difference so it would be the difference first then twice that!

OpenStudy (johnweldon1993):

If it were 8...twice the difference between 8 and 4 i 4 * 2 = 8 not 20....

OpenStudy (johnweldon1993):

2(x - 4) = 20 2x - 8 = 20 2x = 28 x = 14 Recheck....the difference 14 - 4 = 10 Twice that = 20

OpenStudy (ankit042):

No answer is not 8 and even I made the mistake it is 2(x-4) = 20 as @johnweldon1993 said Sorry for the fumble

OpenStudy (johnweldon1993):

Can you see that @apop ?